OpenStudy (anonymous):

tell whether this is a ellipse, circle, parabola, or hyperbola. 4x^2-x^2+4=0

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=4x%5E2-x%5E2%2B4%3D0 it's a parabula ;)

OpenStudy (anonymous):

waitt..i needa know why or how u knw!!

OpenStudy (anonymous):

ok ok, resolve it right?

OpenStudy (anonymous):

put a variable t=x^2

OpenStudy (anonymous):

wait..wher did the "t" come from?

OpenStudy (anonymous):

so you will have 4t^2-t+4=0

OpenStudy (anonymous):

you create the t for help with calculus

OpenStudy (anonymous):

no@malevolence

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

this isnt calculus..this is alg 2, i dont understand what u are doing

OpenStudy (anonymous):

Ummm... \[4x^2-x^2+4=0 \implies 3x^2+4=0 \implies 3x^2=-4 \implies x^2=-4/3 \implies\]\[ x= \pm \frac{2}{\sqrt3}i\]

OpenStudy (anonymous):

i just needa know wat graph it will make

OpenStudy (anonymous):

XD sorry man I just read x^4 my fault!

OpenStudy (anonymous):

decripter seems to knw wat im looking for...malevolence..i jus needa knw why its a parabola

OpenStudy (anonymous):

i didnt mean for u to get upset@malevolence

OpenStudy (anonymous):

I can tell you that if you take the x^2 you will have x^2*(4-1+4/(x^2)) so the graph with bigger value of x is like a parabola, but this is analysis I XD

OpenStudy (anonymous):

okay...i will just go ask sumone else thnks for all of ur help decripter...i kinda got wat ur saying..but ur calc language is strong! hahaha..thnks thou

OpenStudy (anonymous):

XD i'm italian, I have some difficulties to write in english XD, sorry ;)

OpenStudy (anonymous):

naw..ur english is perfect! i didnt even detect a mistake! haha u have a good calc language

OpenStudy (anonymous):

:D thank's man XD

OpenStudy (anonymous):

np! thnks for the help :D